# Emergent chirality and superfluidity of parity-doubled baryons in neutron stars

### v1

@@ III.4 Dirac points
in the Nambu-Gor’kov space for the given mean fields, $\boldsymbol{\Delta}_{a}$ and $\boldsymbol{\delta}$, for the ${\cal S}_{a}1^{-}$ and ${\cal A}1^{+}$ condensates, respectively. The eigenvalues of the Hamiltonain (37) are given by the single-particle energies given in Eqs. (21) and (22). Considering the ${\cal A}1^{+}$ condensate, we find that the Hamiltonian (37) is expressed by
@@
The explicit forms of $U_{m}^{\pm}$ are shown in Eqs. (200) in Appendix D.
@@ IV Conclusion and outlook
We have discussed the emergent chiral superfluids of neutron matter, where the neutron $n(940)$ ($J^{P}=1/2^{+}$) and the chiral partner $n^{\ast}(1535)$ ($1/2^{-}$) are degenerate forming parity doubling, at sufficiently high density. Apart from the usual chiral symmetry, we have introduced the $G_{\mathrm{L}}\times G_{\mathrm{R}}=\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries as proper symmetries in the parity-doublet model since the naive and mirror assigments are naturally included as their subgroups. The emergent chiral symmetries are novel symmetries proposed in this study for the first time. Introducing the four-point interactions for $n$ and $n^{\ast}$, we have classified the Lagrangian in each emergent chiral symmetry and analyzed the Cooper pairings based on the Nambu-Gor’kov formalism. The emergent chiral superfluids are characterized by several different condensates: the emergent-chiral-symmetric and vector condensate (${\cal S}_{a}1^{-}$; $a=0,1,3$) and the emergent-chiral-asymmetric and axialvector condensate (${\cal A}1^{+}$). For the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry, we have obtained (i) either ${\cal S}_{0}1^{-}$ or ${\cal S}_{3}1^{-}$ and (ii) either ${\cal S}_{1}1^{-}$ or ${\cal A}1^{+}$ according to the relative strengths in the interaction terms. For the $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry, one of the ${\cal S}_{0}1^{-}$, ${\cal S}_{1}1^{-}$, ${\cal S}_{3}1^{-}$ and ${\cal A}1^{+}$ condensates is formed. The emergent chiral symmetry as well as the spatial rotational symmetry are broken by these condensates. We have analyzed the patterns of symmetry breaking and the appearance of the NG bosons consisting of six quarks, and found that the emergent chiral superfluids exhibit several topological excitations according to the homotopy groups. We have also shown the existence of the (massless) Dirac fermions at Dirac points in momentum space, indicating an anistropy in the transport phenomena.
@@
In this paper, we have considered condensates of only one of the vector condensations $\boldsymbol{\Delta}_{a}$ ($a=0,1,3$) and $\boldsymbol{\delta}$ for ${\cal S}_{0}1^{-}$, ${\cal S}_{1}1^{-}$, ${\cal S}_{3}1^{-}$ and ${\cal A}1^{+}$ for simplicity. We have further assumed for each case a condensation of only one spatial component of each three-dimensional vector condensation, $\boldsymbol{\Delta}_{a}=(0,0,\Delta_{a})$ or $\boldsymbol{\delta}=(0,0,\delta)$, which is not the most general. This situation corresponds to the polar phase in spinor BECs, in which case depending on couplings there is also another phase, the so-called ferromagnetic phase where the $\mathrm{SO}(3)$ rotational symmetry is completely broken spontaneously with magnetizations [84]. Investigating a possibility of such a ferromagnetic phase with a possible origin of magnetic fields in neutron stars is one of future directions. It is also interesting to study whether a simultaneous condensation of ${\cal S}_{0}1^{-}$ and ${\cal S}_{3}1^{-}$, and that of ${\cal S}_{1}1^{-}$ and ${\cal A}1^{+}$ can occur in general. In the ${\cal S}_{0}1^{-}$ or ${\cal S}_{3}1^{-}$ phase in Fig. 5, the ground states of the condensates ${\cal S}_{0}1^{-}$ and ${\cal S}_{3}1^{-}$ are energetically degenerated, while in the ${\cal S}_{1}1^{-}$ or ${\cal A}1^{+}$ phase, those of ${\cal S}_{1}1^{-}$ and ${\cal A}1^{+}$ are degenerated. Thus, there may exist domain walls interpolating these two condensates in each phase.
@@
As discussed in Sec. III.3, there exists a rich variety of topological objects in dynamical symmetry breakings in our superfluids, among which a paricularly important classs is given by superfluid vortices; Since neutron stars are rapidly rotating, there appear a large number of superfluid vortices along the rotation axis and they may form an Abrikosov’s lattice. In our case, a singly quantized vortex may be split into a pair of half-quantized Alice strings [91, 92]. Such superfluid vortices may be important for pulsar glitches, as was proposed for ${}^{3}P_{2}$-${}^{1}S_{0}$ interface [53]. If our superfluids are connected to ${}^{3}P_{2}$ neutron superfluids as illustrated in Fig. 1, half-quantized Alice strings in the former may be connected to half-quantized non-Abelian vortices [37, 49, 50, 47] in the latter.
@@
Anisotropic transport of Dirac fermions may have impact on neutron star cooling. It is an interesting subject to study the interaction between neutrinos and Dirac fermions. There we will also need to consider the NG bosons as well and study the coupled system among neutrinos, Dirac fermions and NG bosons. The Berry phase around the Dirac cone will lead to a nontrivial topological charge showing properties of topological materials.
@@
Studying the NG bosons consisting of six quarks in the neutron emergent chiral superfluids may be an interesting subject for multiquark systems in the light of exotic hadrons by the quark number counting. Similar situations can be considered in the neutron ${}^{1}S_{0}$ and ${}^{3}P_{2}$ superfluids. Revealing the properties of these NG bosons, such as dispersion relations, interactions with medium and so on, at high density may give us a hint to understand multiquark systems in a way different from the conventional hadron spectroscopy in vacuum. Investigating contributions of these NG bosons to neutron star cooling is also an important future problem.
@@
Our results in this paper are obtained exclusively with the specific part of the entire Lagrangian, $\mathcal{L}_{\rm com}$. The given classifications of symmetries and operators both in the naive and mirror assignments may suggest further richness of the phase structure in a more general situation. It is of particular interest to explore the full theory to possibly pin down any characteristics as distinct properties of the given chirality-assignment in high-density QCD.
Acknowledgements
@@ Appendix B Lagrangians in naive and mirror assignments
In Sec. II, we have introduced the Lagrangian with four-point interaction ${\cal L}_{\mathrm{com}}$, which is commonly included both in the naive and mirror assignments. In this appendix, we show the Lagrangian belonging to each of the naive assignment and the mirror assignment. The Lagrangian in the naive assignment is given by
@@ C.2 $\mathrm{SU}(N)$ symmetry
We summarise the Fierz identities from the $\mathrm{SU}(N)$ generators. The generators of $\mathrm{SU}(N)$ group are given by $\tau_{a}$ ($a=1,2,\dots,N^{2}-1$), which are normalized as $\mathrm{tr}(\tau_{a}\tau_{b})=2\delta_{ab}$. $\mathbbm{1}$ is the unit matrix with $N\times N$ dimension, and $\tau_{0}=\sqrt{2/N}\mathbbm{1}$. The Fierz identities for the particle-antiparticle channels are gven by

### v2

@@ III.4 Dirac points
in the Nambu-Gor’kov space for the given mean fields, $\boldsymbol{\Delta}_{a}$ and $\boldsymbol{\delta}$, for the ${\cal S}_{a}1^{-}$ and ${\cal A}1^{+}$ condensates, respectively. The eigenvalues of the Hamiltonian (37) are given by the single-particle energies given in Eqs. (21) and (22). Considering the ${\cal A}1^{+}$ condensate, we find that the Hamiltonian (37) is expressed by
@@
The explicit forms of $U_{m}^{\pm}$ are shown in Eq. (200) in Appendix D.
@@ IV Conclusion and outlook
We have discussed the emergent chiral superfluids of neutron matter, where the neutron $n(940)$ (mass 940 MeV, spin-parity $J^{P}=1/2^{+}$) and the chiral partner $n^{\ast}(1535)$ (mass 1535 MeV, spin-parity $J^{P}=1/2^{-}$) are degenerate forming parity doubling at sufficiently high density. Apart from the usual chiral symmetry, we have introduced the $G_{\mathrm{L}}\times G_{\mathrm{R}}=\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ and $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetries as proper symmetries in the parity-doublet model since the naive and mirror assignments are naturally included as their subgroups. The emergent chiral symmetries are novel symmetries proposed in this study for the first time. Introducing the four-point interactions for $n$ and $n^{\ast}$, we have classified the Lagrangian in each emergent chiral symmetry and analyzed the Cooper pairings based on the Nambu-Gor’kov formalism. The emergent chiral superfluids are characterized by several different condensates: the emergent-chiral-symmetric and vector condensate (${\cal S}_{a}1^{-}$; $a=0,1,3$) and the emergent-chiral-asymmetric and axialvector condensate (${\cal A}1^{+}$). For the $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ emergent chiral symmetry, we have obtained (i) either ${\cal S}_{0}1^{-}$ or ${\cal S}_{3}1^{-}$ and (ii) either ${\cal S}_{1}1^{-}$ or ${\cal A}1^{+}$ according to the relative strengths in the interaction terms. For the $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$ emergent chiral symmetry, one of the ${\cal S}_{0}1^{-}$, ${\cal S}_{1}1^{-}$, ${\cal S}_{3}1^{-}$ and ${\cal A}1^{+}$ condensates is formed. The emergent chiral symmetry as well as the spatial rotational symmetry are broken by these condensates. We have analyzed the patterns of symmetry breaking and the appearance of the NG bosons consisting of six quarks, and found that the emergent chiral superfluids exhibit several topological excitations according to the homotopy groups. We have also shown the existence of the (massless) Dirac fermions at Dirac points in momentum space, indicating an anisotropy in the transport phenomena.
@@
In this paper, we have considered condensates of only one of the vector condensations $\boldsymbol{\Delta}_{a}$ ($a=0,1,3$) and $\boldsymbol{\delta}$ for ${\cal S}_{0}1^{-}$, ${\cal S}_{1}1^{-}$, ${\cal S}_{3}1^{-}$ and ${\cal A}1^{+}$ for simplicity. We have further assumed for each case a condensation of only one spatial component of each three-dimensional vector condensation, $\boldsymbol{\Delta}_{a}=(0,0,\Delta_{a})$ or $\boldsymbol{\delta}=(0,0,\delta)$, which is not the most general. This situation corresponds to the polar phase in spinor BECs, in which case depending on couplings there is also another phase, the so-called ferromagnetic phase where the $\mathrm{SO}(3)$ rotational symmetry is completely broken spontaneously with magnetizations [87]. Investigating a possibility of such a ferromagnetic phase with a possible origin of magnetic fields in neutron stars is one of future directions. It is also interesting to study whether a simultaneous condensation of ${\cal S}_{0}1^{-}$ and ${\cal S}_{3}1^{-}$ and that of ${\cal S}_{1}1^{-}$ and ${\cal A}1^{+}$ can occur in general. In the ${\cal S}_{0}1^{-}$ or ${\cal S}_{3}1^{-}$ phase in Fig. 5, the ground states of the condensates ${\cal S}_{0}1^{-}$ and ${\cal S}_{3}1^{-}$ are energetically degenerated, while in the ${\cal S}_{1}1^{-}$ or ${\cal A}1^{+}$ phase those of ${\cal S}_{1}1^{-}$ and ${\cal A}1^{+}$ are degenerated. Thus, there may exist domain walls interpolating these two condensates in each phase.
@@
As discussed in Sec. III.3, there exists a rich variety of topological objects in dynamical symmetry breakings in our superfluids, among which a particularly important class is given by superfluid vortices; since neutron stars are rapidly rotating, there appear a large number of superfluid vortices along the rotation axis and they may form an Abrikosov’s lattice. In our case, a singly quantized vortex may be split into a pair of half-quantized Alice strings [96, 97]. Such superfluid vortices may be important for pulsar glitches, as was proposed for ${}^{3}P_{2}$-${}^{1}S_{0}$ interface [54]. If our superfluids are connected to ${}^{3}P_{2}$ neutron superfluids as illustrated in Fig. 1, half-quantized Alice strings in the former may be connected to half-quantized non-Abelian vortices [38, 50, 51, 48] in the latter.
@@
Anisotropic transport of Dirac fermions may have an impact on neutron star cooling. It is an interesting subject to study the interaction between neutrinos and Dirac fermions. There we will also need to consider the NG bosons and study the coupled system among neutrinos, Dirac fermions and NG bosons. The Berry phase around the Dirac cone will lead to a nontrivial topological charge showing properties of topological materials.
@@
Studying the NG bosons consisting of six quarks in the emergent chiral superfluids may be an interesting subject for multiquark systems in the light of exotic hadrons by the quark number counting. Similar situations can be considered in the neutron ${}^{1}S_{0}$ and ${}^{3}P_{2}$ superfluids. Revealing the properties of these NG bosons, such as dispersion relations, interactions with medium and so on, at high density may give us a hint to understand multiquark systems in a way different from the conventional hadron spectroscopy in vacuum. Investigating contributions of these NG bosons to neutron star cooling is also an important future problem.
@@
Our results in this paper are obtained exclusively with the specific part of the entire Lagrangian, $\mathcal{L}_{\rm com}$. The given classifications of symmetries and operators both in the naive and mirror assignments may suggest further richness of the phase structure in a more general situation. It is of particular interest to explore the full theory to possibly pin down any characteristics as distinct properties of the given chirality assignment in high-density QCD. Besides, it may also be intriguing to clarify a potential link of the emergent chiral symmetry to the chiral spin symmetry proposed for the chiral structure of hadrons in hot and dense matter [100, 101, 102, 103], as well as to the quarkyonic phase allowing the hadronic degrees freedom near the Fermi surface of quark matter [104]. Those open problems are left for the future studies.
Acknowledgments
@@ Appendix B Lagrangians in naive and mirror assignments
In Sec. II, we have introduced the Lagrangian with four-point interaction ${\cal L}_{\mathrm{com}}$, which is commonly included both in the naive and mirror assignments. In this Appendix, we show the Lagrangian belonging to each of the naive assignments and the mirror assignments. The Lagrangian in the naive assignment is given by
@@ C.2 $\mathrm{SU}(N)$ symmetry
We summarise the Fierz identities from the $\mathrm{SU}(N)$ generators. The generators of $\mathrm{SU}(N)$ group are given by $\tau_{a}$ ($a=1,2,\dots,N^{2}-1$), which are normalized as $\mathrm{tr}(\tau_{a}\tau_{b})=2\delta_{ab}$. $\mathbbm{1}$ is the unit matrix with $N\times N$ dimension, and $\tau_{0}=\sqrt{2/N}\mathbbm{1}$. The Fierz identities for the particle-antiparticle channels are given by